A balloon is filled with 11 L of helium gas at 1.33 atm. What is the pressure of the balloon if it rises to an altitude where its volume is now 188 L?
1
Expert's answer
2020-06-29T07:44:23-0400

Solution:

This process is isothermal, so it obeys the Boyle-Marriott law.

p1V1 = p2V2 , p2 = p1V1/V2 = 1.33*1.01*10-5Pa*1.1*10-2m3/0.188m3=7860 Pa.

answer: 7860 Pa.